Over the period of review, the very small repetitive events that commenced on the 23 January began increasing in frequency from 19:00 on the 24 January to peak at 5 events per minute. The events range in frequency from two to five hertz and are too small to be locatable. A slight increase in the amplitudes of these events were observed about 06:00 on the 25 January which was enough to produce a recognisable signal on the seismometer at Long Ground. There were several other seismic signals recorded on the Gages and Long ground seismometers that are consistent with small rockfalls and eruptions, two of which produced ash emissions reported at 10:30 and 15:50 on the 25 January.
EDM measurements were carried out on the Galways - Chance's Peak - O'Garra's triangle today. The change in line lengths are less than 5mm, which is within the error of the method. The electronic tilt station at Long Ground continue to show no changes.
Excellent views of the crater was obtained very early this morning and later this evening from the helicopter. Active dome growth continue to be most pronounced in the northern parts of the new dome. New material have built up to a steep angle in the north and continue to collapse into the moat area producing rock falls and associated ash clouds. Spillage of material from the recently collapsed spine to the south continue to deposit debris in the moat area to the south and steaming continues from several areas of the new dome. Strong glowing of the active northern part of the new dome is clearly visible from several areas on the eastern flanks of the volcano and continue to attract nightly sightseers.